# The 4th Dimension for Scaling Model Size

### v1

@@ Abstract
Scaling the size of large language models typically involves 3 dimensions: depth, width, and the number of parameters. In this work, we explore the 4th dimension, virtual logical depth (VLD), which allows increasing the effective algorithmic depth without changing the overall parameter count by reusing parameters within the model. Although parameter reuse itself is not new, its intriguing potential and characteristics in model scaling have not been thoroughly studied. We carefully design controlled experiments and have the following key discoveries on VLD scaling:
  1. 1.
    VLD scaling forces the knowledge capacity of the model to stay almost constant, though with some non-significant variations.
  2. 2.
    VLD scaling enables the reasoning capability to be significantly improved, if the scaling method is properly implemented.
  3. 3.
    The number of parameters is proportional to knowledge capacity, but not reasoning capability. Under certain conditions, it is not necessary to increase parameter count to improve reasoning.
  4. 4.
    The above observations hold for various model configurations and are likely to be generally true under the scope of our experiments.
These findings not only provide useful insights on the future model scaling strategies, but also introduces a even deeper question: Do we really need a lot of parameters to get a really intelligent model? We believe there are many unknown dynamics inside model scaling that needs exploration, so we fully open-source our models and code (https://vldscaling.ngrok.io) to ensure reproducibility and facilitate future explorations.
@@ 1 Introduction
Scaling up large language models (LLMs) is widely recognized as a key factor in enhancing their general capabilities (Kaplan et al., 2020; Hoffmann et al., 2022; Brown et al., 2020; Wei et al., 2022). The general concept of model scaling several directions, including model size scaling, data scaling, compute scaling (Wei et al., 2022), and the recent inference scaling (Snell et al., 2024; Jaech et al., 2024). The model size scaling direction typically occurs along three dimensions: depth, width, and the total number of parameters  [^1] Since depth and width directly influence the number of parameters, these three dimensions are not entirely independent. However, the total parameter count can indeed be adjusted independently by modifying the inner dimension of the MLP layer within a transformer block.. Contemporary LLMs often contain an extremely large number of parameters, sometimes exceeding 100 billion (Wei et al., 2022). While such models can store vast amounts of knowledge, this does not necessarily equate to strong reasoning capabilities. In contrast, the human brain exhibits the opposite characteristics. Human memory is relatively limited—hence the invention of tools like pen and paper to compensate for this weakness—yet humans excel at reasoning (Chollet, 2019; Xu & Poo, 2023; Wang & Li, 2024). This cognitive ability is largely supported by a specialized reasoning core located in the prefrontal cortex (Raichle, 2009).
This fundamental difference in the strengths and weaknesses of language models and the human brain raises critical questions about the efficacy of the current scaling paradigm in achieving super-intelligence. Ideally, a truly intelligent system would consist of a powerful reasoning core complemented by a selective yet essential memory, with additional domain-specific knowledge retrieved externally as needed (Kumaran et al., 2016).
This insight motivates us to explore new scaling dimensions that intentionally limits the knowledge capacity to force it towards improving reasoning capabilities. We find that simple weight-sharing within the standard transformer architecture performs surprisingly well. This introduces the core concept of this manuscript: virtual logical depth (VLD). VLD is achieved by repeating transformer layers while sharing weights across these repetitions. Although not entirely new, the intriguing properties of VLD in model scaling is not thoroughly explored. Previous studies (Lan et al., 2020; Dehghani et al., 2019; Bai et al., 2019; Hay & Wolf, 2024; Bhojanapalli et al., 2021; Reid et al., 2021; Takase & Kiyono, 2021; Liu et al., 2023) have discussed parameter reuse or layer repetition in different contexts, but their primary focus was not on understanding its dynamics in scaling. The looped transformers (Saunshi et al., 2025) is a pioneering work that studies the reasoning capability among the related work, but we focus more on the model scaling strategy and provide a set of more comprehensive experiments and new conclusions on the VLD scaling law.
@@ 2 Virtual Logical Depth
VLD is the effective algorithmic depth of the model minus the number of base layers. See Figure 2 for a brief illustration. In this example, there are 3 base layers and 3 VLD layers. We reuse the base layer parameters in different patterns (Takase & Kiyono, 2021) and produce different types of VLD layers. While many previous work on parameter reuse (Reid et al., 2021; Liu et al., 2023) attempted to modify the standard transformer architecture, our focus is not a new architecture but a new scaling dimension under a general architecture. Therefore, we keep the standard transformer architecture and only repeat its layers while sharing their weights. We consider three patterns of parameter reuse: sequence, cycle and inverse cycle, as illustrated in Figure 2 (b-d). In the sequence mode, the neighboring layers share the same parameter. In the cycle mode, a number of layers are grouped into a large block and then the block is repeated. In the inverse cycle mode, a number of layers are grouped into a large block but the block is reversed when it is repeated. We define the core question that we want to explore as follows:
@@
The above question requires us to develop reasonable approaches to measure the knowledge capacity and reasoning capability. Our high-level guideline is that such kind of measurement should be done with rigorously controlled experiments like in Allen-Zhu & Li (2024). We got inspiration from Allen-Zhu & Li (2024) and develop the measurement approaches as explained in Section 2.1 and Section 2.2.
@@ 2.1 Measurement of Knowledge Capacity
First notice that this experiment only relies on a training set but not a validation set, because the objective is to let the LLM remember the training set as good as possible. Once the training is converged, let the softmax probability output at the $j^{th}$ input for the $i^{th}$ value of $X$ is $p_{j}(x_{i})$. To be clear, $p_{j}(x_{i})$ is the probability that the $(j+1)^{th}$ number in the sequence is $x_{i}$, predicted by the LLM. The LLM is expected to output a higher probability if that number is really $x_{i}$ and lower probability otherwise. If the LLM remembers everything and has absolute confidence, then $p_{j}(x_{i})$ is $1$ for the correct $x_{i}$ and is 00 otherwise. But in reality the LLM has limited knowledge capacity. We define quantity $H_{2}$ as:
@@ 2.2 Measurement of Reasoning Capability
Developing a rigorous measurement of reasoning capability is harder than measuring the knowledge capacity. Reasoning capability does not have a clear unit like information entropy bits as we have in knowledge capacity. Therefore, we adopt a widely used empirical metric as a measurement, which is the accuracy of solving mathematical problems. However, we cannot directly use the currently available public training and validation sets because there might be amination issues, where the validation data is somehow presented in the training data to make the model overfit (Xu et al., 2024).

### v3

# Beyond Parameters: Exploring Virtual Logic Depth for Scaling Laws

@@ Abstract
Scaling the size of large language models typically involves 3 dimensions: depth, width, and the number of parameters. In this work, we explore a 4th dimension: virtual logical depth (VLD), which allows increasing the effective algorithmic depth without changing the overall parameter count by reusing parameters within the model. While parameter reuse is not new, its role in scaling dynamics has remained underexplored. Unlike currently trending test-time methods, which mainly scale in token-wise, VLD alters the internal computation graph scaling during training, inference, or combination. We carefully design controlled experiments and have the following key insights on VLD scaling: 1. Knowledge capacity vs. parameters. At a fixed parameter count, VLD leaves knowledge capacity nearly unchanged (with only minor variance), while across models knowledge capacity scales with the number of parameters; 2. Reasoning vs. reuse. Properly implemented VLD substantially improves reasoning ability without increasing parameter count, decoupling reasoning from sheer model size. This provides a new possibility for scaling besides the current token-wise test-time scaling used by most reasoning models. 3. Robustness and generality. The trend of improved reasoning persist across architectures and configurations (e.g., different reuse schedules and step counts), indicating that VLD captures a general scaling behavior. These findings not only provide useful insights into the future model scaling strategies, but also introduce an even deeper question: Does super intelligence necessarily require ever-larger models, or could it have some trade-offs by re-using parameters and increasing virtual logic depth? We believe that there are many unknown dynamics within the model scaling that need exploration. Codes are available at https://anonymous.4open.science/r/virtual_logical_depth-8024/.
@@ 1 Introduction
Scaling up large language models (LLMs) has long been regarded as a primary driver of their rapid progress (Kaplan et al., 2020; Hoffmann et al., 2022; Brown et al., 2020; Wei et al., 2022). Scaling typically proceeds along multiple axes, including model size, data, compute (Wei et al., 2022), and more recently inference-time scaling (Snell et al., 2024; Jaech et al., 2024). The model size scaling direction typically occurs along three dimensions: depth, width, and the total number of parameters  [^1] Since depth and width directly influence the number of parameters, these three dimensions are not entirely independent. However, the total parameter count can indeed be adjusted independently by modifying the inner dimension of the MLP layer within a transformer block.. Contemporary LLMs often exceed 100B parameters (Wei et al., 2022), enabling them to store vast knowledge. However, parameter growth does not straightforwardly translate into stronger reasoning capabilities.
In contrast, human cognition demonstrates the opposite profile. Memory is relatively limited—hence the invention of external storage such as writing—yet reasoning and abstract problem solving are remarkably strong (Chollet, 2019; Xu & Poo, 2023; Wang & Li, 2024), largely supported by specialized circuits in the prefrontal cortex (Raichle, 2009). This asymmetry highlights a fundamental tension: while current scaling paradigms expand memory-like capacity, they do not necessarily improve reasoning, which is arguably central to achieving super-intelligence. Ideally, an intelligent system would combine (i) a powerful reasoning core, (ii) a compact but essential internal memory, and (iii) external retrieval for domain knowledge (Kumaran et al., 2016).
Motivated by this perspective, we introduce a new scaling dimension that deliberately limits knowledge capacity in order to encourage improvements in reasoning. Specifically, we revisit weight sharing in the transformer architecture and show that repeating layers with tied parameters yield surprisingly strong benefits. We term the resulting phenomenon virtual logical depth (VLD). While related techniques—such as parameter reuse, layer repetition, and equilibrium models (Lan et al., 2020; Dehghani et al., 2019; Bai et al., 2019; Hay & Wolf, 2024; Bhojanapalli et al., 2021; Reid et al., 2021; Takase & Kiyono, 2021; Liu et al., 2023)—have been explored in other contexts, their scaling dynamics have remained underexplored. A closely related effort, looped transformers (Saunshi et al., 2025), studies reasoning capacity, but our focus is on scaling laws: we conduct systematic experiments and uncover new properties of VLD as a scaling strategy.
@@
2 Related Work
Parameter Sharing and Layer Reuse.
Parameter reuse has garnered considerable attention in recent research. (Reid et al., 2021) introduced Subformer, which reused 50% of parameters for compression purposes, while (Liu et al., 2023) presented ESP (Efficient Shared Parameterization) that reused central tensor matrices. However, both introduced atypical elements to transformer architectures, complicating direct comparison with standard models. (Bhojanapalli et al., 2021) developed Reuse Transformers, exploring attention mechanism reuse while sharing less than 10% of attention heads. (Takase & Kiyono, 2021) proposed three distinct layer-sharing patterns but conducted experiments only on small networks without scaling analysis. More recently, (Hay & Wolf, 2024) proposed Dynamic Layer Tying, employing reinforcement learning to identify optimal layer-sharing configurations. In parallel, several studies have explored Looped Transformers (Fan et al., 2024; Saunshi et al., 2025; Yang et al., 2023), which iteratively apply transformer layers to enhance model capabilities without proportionally increasing parameters. Our work builds upon these foundations by systematically analyzing VLD scaling dynamics and its distinct impact on reasoning capability versus knowledge capacity.
Measuring Reasoning and Knowledge Capacity.
Quantitative assessment of reasoning and knowledge remains challenging. For reasoning capabilities, (Ye et al., 2024) employed synthetic GSM8K benchmarks to isolate inherent reasoning abilities from memorized solutions. Regarding knowledge capacity, (Allen-Zhu & Li, 2024) demonstrated that LLMs store approximately 2 bits of knowledge per parameter, providing a quantitative metric for model memory. (Carlini et al., 2022) established methodologies for quantifying memorization in LLMs, offering insights into information retention mechanisms. Understanding how large language models (LLMs) store and retrieve information is central to evaluating their memory mechanisms. (Liang et al., 2024) introduced the SAGE framework, which enhances LLM agents with reflective and memory-augmented abilities. (Xu et al., 2025) propose A-Mem, an agentic memory framework designed to enhance the long-term reasoning and adaptability of LLM agents.
@@ 3 Virtual Logical Depth
We define Virtual Logical Depth (VLD) as the effective algorithmic depth of a model minus the number of base layers. Figure 2 provides an illustration: here, 3 base layers are reused to form 3 VLD layers. Instead of introducing a new architecture, our goal is to establish a new scaling dimension within the standard transformer framework. Concretely, we retain the transformer architecture but repeat its layers with shared parameters, following the parameter reuse insights of Takase & Kiyono (2021).
We explore three reuse patterns—sequence, cycle, and inverse cycle—as shown in Figure 2 (b–d). In the sequence mode, adjacent layers share parameters. In the cycle mode, several layers form a block that is repeated. In the inverse cycle mode, the same block is repeated but in reversed order. Detailed implementation and intuition for each strategy are described in Appendix D.
@@
The above question requires us to develop reasonable approaches to measure the knowledge capacity and reasoning capability. Our high-level guideline is that such kind of measurement should be done with rigorously controlled experiments like in (Allen-Zhu & Li, 2024). We got inspiration from (Allen-Zhu & Li, 2024) and develop the measurement approaches as explained in Section 3.1 and Section 3.2.
@@ 3.1 Measurement of Knowledge Capacity
First notice that this experiment only relies on a training set but not a validation set, because the objective is to let the LLM remember the training set as good as possible. Once the training is converged, let the softmax probability output at the $j^{th}$ input for the $i^{th}$ value of $X$ is $p_{j}(x_{i})$. To be clear, $p_{j}(x_{i})$ is the probability that the $(j+1)^{th}$ number in the sequence is $x_{i}$, predicted by the LLM. The LLM is expected to output a higher probability if that number is really $x_{i}$ and lower probability otherwise. If the LLM remembers everything and has absolute confidence, then $p_{j}(x_{i})$ is $1$ for the correct $x_{i}$ and is 0 otherwise. But in reality the LLM has limited knowledge capacity. We define quantity $H_{2}$ as:
@@ 3.2 Measurement of Reasoning Capability
Developing a rigorous measurement of reasoning capability is harder than measuring the knowledge capacity. Reasoning capability does not have a clear unit like information entropy bits as we have in knowledge capacity. Therefore, we adopt widely used empirical metrics as measurements, including accuracy on mathematical problems, scientific reasoning tasks, and code generation benchmarks. To ensure comprehensive evaluation, we employ both synthetic and real-world datasets to assess VLD’s impact on reasoning capabilities.
3.2.1 Synthetic Data